OpenStudy (anonymous):

What is the simplified form of x plus 2 over x squared minus 3x minus 10 • x minus 3 over x squared plus x minus 12 ? 1 over the quantity x minus 3 times the quantity x plus 4 1 over the quantity x minus 3 times the quantity x plus 2 1 over the quantity x plus 4 times the quantity x minus 5 1 over the quantity x plus 2 times the quantity x minus 5

OpenStudy (anonymous):

\[\frac{ x + 2 }{ x^2 - 3x - 10 } \times \frac{ x - 3 }{ x^2 + x -12 }\]

OpenStudy (anonymous):

@iGreen do i flip the 2nd fraction then horizontally multiply ?

OpenStudy (igreen):

No, that was for division..you can horizontally multiply now.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

quick question when we horizontally multipl would the numerator be x^2 + 2 -3 or x+ 2 x- 3?

OpenStudy (igreen):

None.. \(\sf (x+2)(x-3)\) x * x = ? x * -3 = ? 2 * x = ? 2 * -3 = ?

OpenStudy (anonymous):

x^2 -3x 2 -6

OpenStudy (igreen):

Actually, first we cancel the common factors.

OpenStudy (anonymous):

for the denominator ?

OpenStudy (igreen):

For the 2nd fraction

OpenStudy (igreen):

1st*

OpenStudy (igreen):

Factor \(\sf x^2 - 3x - 10\) from the numerator..tell me what you get.

OpenStudy (igreen):

Can you do that?

OpenStudy (anonymous):

is it x/ x^2 - 3x -5

OpenStudy (igreen):

No..

OpenStudy (igreen):

Can you factor \(\sf x^2 - 3x - 10\)?

OpenStudy (anonymous):

oh yeah give me sec

OpenStudy (igreen):

Okay

OpenStudy (anonymous):

(x + 2) (x - 5)

OpenStudy (igreen):

Yes, so we have: \(\sf \dfrac{x+2}{(x-5)(x+2)} \times \dfrac{x-3}{x^2+x-12}\) The x + 2's cancel out, giving us: \(\sf \dfrac{\cancel{x+2}}{(x-5)(\cancel{x+2})} \times \dfrac{x-3}{x^2+x-12}\) \(\sf \dfrac{1}{(x-5)} \times \dfrac{x-3}{x^2+x-12}\)

OpenStudy (igreen):

Now we multiply horizontally..but first factor \(\sf x^2 + x - 12\).

OpenStudy (anonymous):

(x - 3) (x + 4)

OpenStudy (anonymous):

but it would be \[\frac{ 1 }{ (x - 5) } \times \frac{ 1 } { ( x + 4)}\]

OpenStudy (igreen):

Yes..which gives us: \(\sf \dfrac{1}{(x-5)(x+4)}\)

OpenStudy (igreen):

@shinebrightlikeadimon

OpenStudy (anonymous):

okay thanks so much for some reason i lost connection thats why im replying late